Can you give me the formula and help me to answer the follwoing question: Find the following sum
35-36-37-38....145-145-150

Answer by checkley77(12844)

Sum=n/2(a+L)
n=the number of values.
a=the first value
L=the last value.
Sum=(150-38)/2*(35+150)
Sum=112/2*185
Sum=56*185=10,360 ans.